# Locally similar distances and equality of the induced intrinsic distances

### v1

@@ Abstract
Let $X$ be a set and $d_{1},d_{2}$ be two distances on $X$. We say that $d_{1}$ and $d_{2}$ are locally similar and write $d_{1}\cong d_{2}$ if $d_{1}$ and $d_{2}$ are topologically equivalent and for every $a$ in $X$,
@@
We prove that if $d_{1}\cong d_{2}$, then the intrinsic distances induced by $d_{1}$ and $d_{2}$ are equal. We also provide some sufficient conditions for $d_{1}\cong d_{2}$ and consider several examples related to reproducing kernel Hilbert spaces.
@@ 1 Introduction
The concept of distance is simple, but it plays a fundamental role in mathematics. See [11, 8, 12, 10] for many examples and applications of distances. There are many different notions of “equivalent distances”, including the following ones:
@@
Every distance $d$ induces the intrinsic distance $d^{\ast}$, see [8, Chapter 2] or [12, Chapter 1]. Obviously, the binary relation $d_{1}\stackrel{{\scriptstyle\operatorname{intr}}}{{\sim}}d_{2}$ (“intrinsic equivalence”), which we define by $d_{1}^{\ast}=d_{2}^{\ast}$, is an equivalence relation between the distances on a fixed domain. A natural problem is to obtain criteria for $d_{1}^{\ast}=d_{2}^{\ast}$. It is easy to see that $d_{1}\stackrel{{\scriptstyle\operatorname{Lip}}}{{\sim}}d_{2}$ implies $d_{1}^{\ast}\stackrel{{\scriptstyle\operatorname{Lip}}}{{\sim}}d_{2}^{\ast}$ (Proposition 2.8), but it does not imply $d_{1}^{\ast}=d_{2}^{\ast}$ (Example 2.13).
@@
Our main result (Theorem 6.6) states that the relation $d_{1}\cong d_{2}$ is sufficient for $d_{1}^{\ast}=d_{2}^{\ast}$. We show that the relation $\cong$ is useful to compute $d^{\ast}$ in various examples. Nevertheless, $d_{1}\cong d_{2}$ is not necessary for $d_{1}^{\ast}=d_{2}^{\ast}$; see Examples 2.10, 2.11, and 3.9.
By definition, locally similar distances are topologically equivalent. We show that $\cong$ is independent of $\stackrel{{\scriptstyle\operatorname{Lip}}}{{\sim}}$, and $\stackrel{{\scriptstyle\operatorname{u}}}{{\sim}}$: the biLipschitz equivalence does not imply the local similarity (Example 3.6), and the local similarity does not imply even the uniform equivalence (Example 3.8). See Figure 1.
$\cong$ $\stackrel{{\scriptstyle\operatorname{intr}}}{{\sim}}$ $\stackrel{{\scriptstyle\operatorname{Lip}}}{{\sim}}$ $\stackrel{{\scriptstyle\operatorname{u}}}{{\sim}}$ $\stackrel{{\scriptstyle\operatorname{top}}}{{\sim}}$ Theorem 6.6Example 3.6Example 3.8Example 2.11Example 2.13
Figure 1: Some logic relations between different equivalencies of distances.
Several authors have studied the situation when $d$ is an intrinsic distance, i.e., $d^{\ast}=d$ (see [12, Section 1.3], [11, Chapter 1, parts A, B], [8, Section 2.3]). The constructions of this paper are useful in a more general case where $d^{\ast}\cong d$.
It is natural to relate this work with Riemannian metrics; we hope to do it in the future. In this paper, we avoid Riemannian metrics to show the power of more elementary tools.
@@
In Section 2, we recall the definition of the intrinsic distance associated with a given distance; we also provide examples which are used throughout the paper. In Section 3, we introduce the concept of locally similar distances and show that it is not comparable with $\stackrel{{\scriptstyle\operatorname{Lip}}}{{\sim}}$ or $\stackrel{{\scriptstyle\operatorname{u}}}{{\sim}}$. In Section 4, we study two other concepts (strong local similarity and infinitesimal similarity) that are slightly stronger than local similarity. In Section 5, we show that composing a distance $d_{1}$ with a function that satisfies certain properties, results in a locally similar distance to $d_{1}$ ($d_{2}\cong d_{1}$, if $d_{2}=f\circ d_{1}$). Section 6 contains a proof of the main result. Section 7 contains several examples related to the distances induced by kernels (or, equivalently, by reproducing kernel Hilbert spaces).
2 The intrinsic distance induced by a given distance
In this section, we recall the definition of the intrinsic distance $d^{\ast}$ induced by a given distance $d$. It is a standard construction in the theory of length spaces (see, for example [11, Example 1.4 (a)], [8, Section 2.3], [12, Section 1.2.3]). We include it to introduce the necessary intermediate concepts and fix some notation. We also compute $d^{\ast}$ for some examples that are useful in subsequent sections.
$B_{d}(a,\delta)\coloneqq\bigl\{x\in X\colon\ d(x,a)<\delta\bigr\}.$
Furthermore, we denote by $\tau_{d}$ the topology induced by $d$.
Definition 2.1 (continuous paths between two points).
Let $(X,d)$ be a metric space. For each $x,y$ in $X$, define
Definition 2.2 (Riemannian partitions of the unit interval).
We denote by $\mathcal{P}$ the set of all finite subsets of $[0,1]$ that include 0 and $1$.
Given $P$ in $\mathcal{P}$, we enumerate the elements of $P$ in ascending order:
$P=\{t_{0},t_{1},\ldots,t_{m}\},\qquad 0=t_{0}<t_{1}<\cdots<t_{m}=1.$ (2.1)
Lemma 2.4.
$S(d,\gamma,P)\leq S(d,\gamma,Q).$
Proof.
It follows by an iterative application of the triangle inequality. ∎
Definition 2.5 (the length of a path with respect to a distance).
$L(d,\gamma)\coloneqq\sup_{P\in\mathcal{P}}S(d,\gamma,P).$
Definition 2.6 (the intrinsic distance induced by a distance).
$d^{\ast}(x,y)\coloneqq\inf_{\gamma\in\Gamma_{d}(x,y)}L(d,\gamma).$
Following Definition 2.6, if there are not any continuous paths between $x$ and $y$ (i.e. $\Gamma_{d}(x,y)=\emptyset$), then, $d^{\ast}(x,y)=+\infty$.
It is well known that $d^{\ast}$ is a $0,+\infty]-$valued distance on $X$, such that $d^{\ast}(x,y)\geq d(x,y)$ for every $x,y$ in $X$ (see the references cited at the beginning of this section).
Remark 2.7.
In this paper, we define paths as continuous functions with domain $[0,1]$. Similarly, one can consider paths defined on arbitrary closed intervals of the form $[a,b]$. Such generalization does not change $L(d,\gamma)$ or $d^{\ast}(x,y)$, since any path on an arbitrary interval $[a,b]$ can be reparametrized to $[0,1]$.
Proposition 2.8.
Proof.
Let $x,y\in X$ and $\gamma\in\Gamma_{d_{1}}(x,y)$. For every $P$ in $\mathcal{P}$,
$S(d_{2},\gamma,P)\leq C_{2}S(d_{1},\gamma,P)\leq C_{2}L(d_{1},\gamma).$
Example 2.9 (Euclidean distance on $\mathbb{R}^{n}$).
Example 2.10 (discrete distance on $\mathbb{R}$).
Define $d_{0}\colon\mathbb{R}^{2}\to[0,+\infty)$,
$d_{0}(x,y)\coloneqq\begin{cases}0,&x=y;\\ 1,&x\neq y.\end{cases}$
The topology induced by $d_{0}$ is the discrete topology. Therefore, there are no continuous paths between different points of $\mathbb{R}$, and $d_{0}^{\ast}(x,y)=+\infty$ if $x\neq y$. ∎
Example 2.11 (square-root distance on $\mathbb{R}$).
Consider the distance $d_{\operatorname{sqrt}}\colon\mathbb{R}^{2}\to[0,+\infty)$, given by $d_{\operatorname{sqrt}}(x,y)\coloneqq\sqrt{|x-y|}$. The topology induced by $d_{\operatorname{sqrt}}$ is the usual topology in $\mathbb{R}$, since the set of open balls with respect to the Euclidean distance in $\mathbb{R}^{2}$ is the same as the set of open balls with respect to the distance $d_{\operatorname{sqrt}}$.
Let $x,y\in\mathbb{R}$. We will show that if $x\neq y$, then $d_{\operatorname{sqrt}}^{\ast}(x,y)=+\infty$. Suppose $x<y$. Let $\gamma\in\Gamma_{d_{\operatorname{sqrt}}}(x,y)$. Notice that $\gamma$ is continuous in the usual sense. Let $m\in\mathbb{N}$. For every $j$ in $\{0,\ldots,m\}$, we define
$z_{m,j}\coloneqq x+\frac{j}{m}(y-x).$
$t_{m,m}\coloneqq 1,$
$t_{m,0}<t_{m,1}<\cdots<t_{m,m}.$
Taking the supremum over $m$ in $\mathbb{N}$ shows that $L(d_{\operatorname{sqrt}},\gamma)=+\infty$. Since $\gamma$ was an arbitrary path between $x$ and $y$, it follows from Definition 2.6, that $d_{\operatorname{sqrt}}^{\ast}(x,y)=+\infty$ if $x\neq y$. ∎
Although the distances $d_{0}$ and $d_{\operatorname{sqrt}}$ from Examples 2.10 and 2.11, are not topologically equivalent, ther corresponding intrinsic distances are, in fact, equal: $d_{0}^{\ast}=d_{\operatorname{sqrt}}^{\ast}$.
Example 2.12 (pseudologarithmic distance on the halfline).
Let $X\coloneqq(0,+\infty)$. Define $\rho\colon X^{2}\to[0,+\infty)$,
$\rho(x,y)\coloneqq\frac{2|x-y|}{x+y}.$
We are going to prove that
$\rho^{\ast}(x,y)=|\log(x)-\log(y)|.$ (2.2)
Consider the case where $x<y$ (the case $x>y$ is similar). Let $\xi(t)\coloneqq x+t(y-x)$ be the straight-line path between $x$ and $y$, and let $Q_{m}=\{t_{0},\ldots,t_{m}\}$ be an arbitrary partition. For each $k$ in $\{0,1,\ldots,m\}$, we define $z_{k}\coloneqq\xi(t_{k})$. This setup leads to an explicit formula for the polygonal length of $\xi$:
$S(\rho,\xi,Q_{m})=\sum_{k=0}^{m-1}\rho(z_{k},z_{k+1})=\sum_{k=0}^{m-1}\frac{2(z_{k+1}-z_{k})}{z_{k+1}+z_{k}}=\sum_{k=0}^{m-1}\frac{z_{k+1}-z_{k}}{\frac{z_{k+1}+z_{k}}{2}}.$
The resulting sum is a Riemannian sum for the integral of the function $f(t)=\frac{1}{t}$ over the interval $[x,y]$. Therefore,
$L(\rho,\xi)=\lim_{m\to\infty}S(\rho,\xi,Q_{m})=\int_{x}^{y}\frac{\mathrm{d}{}t}{t}=\log(y)-\log(x).$
For an arbitrary $\gamma$ in $\Gamma_{\rho}(x,y)$, we can construct a partition $P_{m}=\{t_{m,0},\ldots,t_{m,m}\}$ with $\gamma(t_{m,k})=x+k(y-x)/m$, as in Example 2.11, and obtain
$L(\rho,\gamma)\geq S(\rho,\gamma,P_{m})\geq\sum_{k=0}^{m-1}\rho\left(x+k\frac{y-x}{m},x+(k+1)\frac{y-x}{m}\right).$
The right-hand side is also a Riemannian sum for the integral $\int_{x}^{y}\mathrm{d}{}t/t$. Passing to the limit as $m\to\infty$ we get $L(\rho,\gamma)\geq\log(y)-\log(x)$. Thereby, we have proven (2.2). ∎
Example 2.13 (pseudologarithmic distance on a segment).
Let $X\coloneqq[1,2]$. Define $\rho\colon X^{2}\to[0,+\infty)$ as
$\rho(x,y)\coloneqq\frac{2|x-y|}{x+y}.$
Example 2.14 (pseudohyperbolic distance on the halfplane).
Let $\mathbb{H}$ be the upper halfplane of the complex plane:
$\mathbb{H}\coloneqq\{z\in\mathbb{C}\colon\ \operatorname{Im}(z)>0\}.$
Define $\rho_{\mathbb{H}}\colon\mathbb{H}^{2}\to[0,+\infty)$,
$\rho_{\mathbb{H}}(ia,ib)=\frac{|a-b|}{a+b}\qquad(a,b>0).$
Indeed, for $z=x+iy$ and $w=u+iv$,
Example 2.15 (pseudohyperbolic distance on the unit disk).
It is well known (see, e.g., [18, Chapter 33] and [21, Section 1.5]) that $2\rho_{\mathbb{H}}^{\ast}$ is the geodesic distance associated with the Poincaré halfplane metric $\mathrm{d}{}x\,\mathrm{d}{}y/y^{2}$, and $2\rho_{\mathbb{D}}^{\ast}$ is the geodesic distance associated with the Poincaré disc metric $\mathrm{d}{}x\,\mathrm{d}{}y/(1-x^{2}-y^{2})^{2}$. Nevertheless, we emphasize that the intrinsic distances $\rho_{\mathbb{H}}^{\ast}$ and $\rho_{\mathbb{D}}^{\ast}$ in these classical examples can be computed without using Riemannian metrics.
@@ 3 Locally similar distances
$\forall\varepsilon>0\qquad\exists\delta>0\qquad\forall x\in B_{d_{2}}(a,\delta)\setminus\{a\}\qquad\left|\frac{d_{1}(x,y)}{d_{2}(x,y)}-1\right|<\varepsilon.$ (3.6)
@@
Example 3.9.
Let $X$ be the following “comb” in $\mathbb{R}^{2}$; see Figure 3:
Furthermore, $d^{\ast}=\rho$. In other words, we claim that $d^{\ast}(a,b)=\rho(a,b)$ for every $a$, $b$ in $X$. For brevity, we illustrate the idea of the proof for the particular case where $a=(0,0)$ and $b=(1/q,y)$, $q\in\mathbb{N}$, $0<y\leq 1$. If $\gamma\colon[0,1]\to X$ is a continuous path joining $a$ with $b$, then, by connectivity arguments, there exists $u$ in $(0,1)$ such that $\gamma(t)=(1/q,0)$. It is easy to see that $L(d,\gamma|_{[0,u]})\geq 1/q$ and $L(d,\gamma|_{[u,1]})\geq y$; here the domains of the paths are $[0,u]$ and $[u,1]$; see Remark 2.7. Therefore, $L(d,\gamma)\geq\rho(a,b)$ for every $\gamma$ in $\Gamma_{d}(a,b)$, and it is easy to conclude that $d^{\ast}(a,b)=\rho(a,b)$. In particular, $d\leq\rho$.
@@ 4 Strongly locally similar and infinitesimally similar distances
In this section, we give two alternative concepts for similarity of distances (Definitions 4.1 and 4.5) that are slightly stronger than $\cong$ from Definition 3.1.
@@
$\lim_{\begin{subarray}{c}x,y\to a\\ x\neq y\end{subarray}}\frac{|d_{2}(x,y)-d_{1}(x,y)|}{d_{1}(x,a)+d_{1}(y,a)}=0.$ (4.13)
@@
1. Symmetry. Suppose that $d_{1}\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}d_{2}$. Let $\tau\coloneqq\tau_{d_{1}}$; by definition of $\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}$, we also have $\tau=\tau_{d_{2}}$. Let $a\in X$ and $\varepsilon>0$. From (4.13), there is $V$ in $\tau(a)$ such that for every $x,y$ in $U$,
$\frac{|d_{1}(x,y)-d_{2}(x,y)|}{d_{1}(x,a)+d_{1}(y,a)}<\frac{\min\{\varepsilon,1\}}{2}.$ (4.14)
@@
$\frac{|d_{2}(x,y)-d_{1}(x,y)|}{d_{2}(x,a)+d_{2}(y,a)}\leq\frac{2\,|d_{2}(x,y)-d_{1}(x,y)|}{d_{1}(x,a)+d_{1}(y,a)}<\varepsilon.$
@@
$\frac{|d_{1}(x,y)-d_{2}(x,y)|}{d_{1}(x,a)+d_{1}(y,a)}<\frac{\min\{\varepsilon,1\}}{2},\quad\frac{|d_{2}(x,y)-d_{1}(x,y)|}{d_{2}(x,a)+d_{2}(y,a)}<\frac{\min\{\varepsilon,1\}}{2},$
@@
$\frac{|d_{2}(x,y)-d_{3}(x,y)|}{d_{2}(x,a)+d_{2}(y,a)}<\frac{\min\{\varepsilon,1\}}{2}.$
@@
$\frac{1}{2}\bigl(d_{2}(x,a)+d_{2}(y,a)\bigr)\leq d_{1}(x,a)+d_{1}(y,a).$
So,
$\frac{|d_{1}(x,y)-d_{3}(x,y)|}{d_{1}(x,a)+d_{1}(y,a)}$ $\leq\frac{|d_{1}(x,y)-d_{2}(x,y)|}{d_{1}(x,a)+d_{1}(y,a)}+\frac{|d_{2}(x,y)-d_{3}(x,y)|}{d_{1}(x,a)+d_{1}(y,a)}$
$\leq\frac{|d_{1}(x,y)-d_{2}(x,y)|}{d_{1}(x,a)+d_{1}(y,a)}+\frac{2\,|d_{2}(x,y)-d_{3}(x,y)|}{d_{2}(x,a)+d_{2}(y,a)}<\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon.\qed$
@@
Let $X$ be a set and $d_{1},d_{2}$ be distances on $X$. If $d_{1}\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}d_{2}$, then $d_{1}\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}d_{2}$. If $d_{1}\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}d_{2}$, then $d_{1}\cong d_{2}$.
@@
1. Suppose that $d_{1}\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}d_{2}$. We use the following upper bound for the expression in (4.13):
@@
As $x,y\to a$ and $x\neq y$, the last expression tends to zero.
2. Suppose that $d_{1}\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}d_{2}$. Let $a\in X$. We suppose that the limit relation (4.13) holds. Fixing $y=a$ we get the following consequence of (4.13):
$\lim_{\begin{subarray}{c}x\to a\\ x\neq a\end{subarray}}\frac{|d_{2}(x,a)-d_{1}(x,a)|}{d_{1}(x,a)}=0.$
@@
For two points belonging to the same halfaxe, we have
@@
$\frac{|d_{2}((u,0),(0,u))-d_{1}((u,0),(0,u))}{d_{1}((u,0),(0,0))+d_{1}((0,u),(0,0))}=\frac{2-\sqrt{2}}{2}.$
@@
Problem 4.9.
Is infinitesimal similarity equivalent to strong local similarity?

### v2

@@ Abstract
Let $X$ be a set and $d_{1},d_{2}$ be two distances on $X$. We say that $d_{1}$ and $d_{2}$ are locally similar and write $d_{1}\cong d_{2}$ if $d_{1}$ and $d_{2}$ are topologically equivalent and, for every $a$ in $X$,
@@
We prove that if $d_{1}\cong d_{2}$, then the intrinsic distances induced by $d_{1}$ and $d_{2}$ coincide. We also provide sufficient conditions for $d_{1}\cong d_{2}$ and consider several examples related to reproducing kernel Hilbert spaces.
@@ 1 Introduction
The concept of distance is simple, but it plays a fundamental role in mathematics. See [11, 8, 12, 10] for many examples and applications of distances. There are many different notions of “equivalent distances” on a fixed set $X$, including the following standard notions:
@@
Every distance $d$ induces the intrinsic distance $d^{\ast}$, see [8, Chapter 2] or [12, Chapter 1]. Obviously, the binary relation $d_{1}\stackrel{{\scriptstyle\operatorname{intr}}}{{\sim}}d_{2}$ (“intrinsic equivalence”), which we define by $d_{1}^{\ast}=d_{2}^{\ast}$, is an equivalence relation on the set of all distances defined on a fixed domain. A natural problem is to obtain criteria for $d_{1}^{\ast}=d_{2}^{\ast}$. It is easy to see that $d_{1}\stackrel{{\scriptstyle\operatorname{Lip}}}{{\sim}}d_{2}$ implies $d_{1}^{\ast}\stackrel{{\scriptstyle\operatorname{Lip}}}{{\sim}}d_{2}^{\ast}$ (Proposition 5.8), but it does not imply $d_{1}^{\ast}=d_{2}^{\ast}$ (Example 5.14).
@@
Our main result (Theorem 6.6) states that the relation $d_{1}\cong d_{2}$ is sufficient for $d_{1}^{\ast}=d_{2}^{\ast}$. We show that the relation $\cong$ is useful to compute $d^{\ast}$ in various examples. Nevertheless, $d_{1}\cong d_{2}$ is not necessary for $d_{1}^{\ast}=d_{2}^{\ast}$; see Examples 5.11 and 6.9.
By definition, locally similar distances are topologically equivalent. We show that $\cong$ is independent of $\stackrel{{\scriptstyle\operatorname{Lip}}}{{\sim}}$, and $\stackrel{{\scriptstyle\operatorname{u}}}{{\sim}}$: the biLipschitz equivalence does not imply the local similarity (Example 2.6), and the local similarity does not imply even the uniform equivalence (Example 2.8). Figure 1 shows the logical relationships among the equivalencies of distances considered in the paper.
$\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}$ $\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}$ $\cong$ $\stackrel{{\scriptstyle\operatorname{intr}}}{{\sim}}$ $\stackrel{{\scriptstyle\operatorname{Lip}}}{{\sim}}$ $\stackrel{{\scriptstyle\operatorname{u}}}{{\sim}}$ $\stackrel{{\scriptstyle\operatorname{top}}}{{\sim}}$ Thm. 6.6Ex. 6.9Pr. 3.9Ex. 3.10Pr. 3.8Ex. 3.11Rem. 3.4(by def.)Ex. 2.6Ex. 2.8Ex. 5.12Ex. 5.14
Figure 1: Some logic relations between different equivalencies of distances. We give references to the corresponding theorems, propositions, remarks, and examples (counter-examples).
Several authors have studied the situation when $d$ is an intrinsic distance, i.e., $d^{\ast}=d$ (see [12, Section 1.3], [11, Chapter 1, parts A, B], [8, Section 2.3]). The constructions of this paper are useful in a more general case, namely when $d^{\ast}\cong d$.
We hope to relate this work to Riemannian metrics in future research. In this paper, we avoid Riemannian metrics to show the power of more elementary tools.
@@
In Section 2, we introduce the concept of locally similar distances and show that it is not comparable with $\stackrel{{\scriptstyle\operatorname{Lip}}}{{\sim}}$ or $\stackrel{{\scriptstyle\operatorname{u}}}{{\sim}}$. In Section 3, we study two other concepts (strong local similarity $\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}$ and infinitesimal similarity $\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}$) that are slightly stronger than local similarity. In Section 4, we show that composing a distance $d_{1}$ with a function that satisfies certain properties yields a distance $d_{2}$ that is locally similar to $d_{1}$ ($d_{2}\cong d_{1}$, if $d_{2}=f\circ d_{1}$). In Section 5, we recall the definition of the intrinsic distance associated with a given distance; we also provide examples which are used throughout the paper. Section 6 contains a proof of the main result. Section 7 contains several examples related to the distances induced by kernels (or, equivalently, by reproducing kernel Hilbert spaces).
@@ 2 Locally similar distances
$B_{d}(a,\delta)\coloneqq\bigl\{x\in X\colon\ d(x,a)<\delta\bigr\}.$
Furthermore, we denote by $\tau_{d}$ the topology induced by $d$.
@@
$\forall\varepsilon>0\qquad\exists\delta>0\qquad\forall x\in B_{d_{2}}(a,\delta)\setminus\{a\}\qquad\left|\frac{d_{1}(x,a)}{d_{2}(x,a)}-1\right|<\varepsilon.$ (2.6)
@@ 3 Strongly locally similar and infinitesimally similar distances
In this section, we give two alternative concepts for local similarity of distances (Definitions 3.1 and 3.6) and show that they are slightly stronger than $\cong$ from Definition 2.1. The concepts introduced in this section are close to definitions from Gromov [11, Definition 1.1] and Lytchak [13, Definition 3.2].
@@
Remark 3.4.
In Example 2.8, the distances $d_{X}$ and $\rho$ are strongly locally similar but not uniformly equivalent.
@@
$\lim_{(x,y)\to(a,a)}\frac{|d_{2}(x,y)-d_{1}(x,y)|}{d_{1}(x,a)+d_{1}(y,a)}=0.$ (3.13)
As usual, in the definition of $\lim_{(x,y)\to(a,a)}$ we assume that $(x,y)\neq(a,a)$. To include the point $(a,a)$, we rewrite the limit relation (3.13) in the following equivalent form:
$\forall\varepsilon>0\quad\exists V\in\tau(a)\quad\forall x,y\in V\qquad|d_{2}(x,y)-d_{1}(x,y)|\leq\varepsilon\,\bigl(d_{1}(x,a)+d_{1}(y,a)\bigr),$ (3.14)
where $\tau$ is the topology induced by $d_{1}$ (or by the topologically equivalent distance $d_{2}$).
@@
1. Symmetry. Suppose that $d_{1}\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}d_{2}$. Let $\tau\coloneqq\tau_{d_{1}}$; by definition of $\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}$, we also have $\tau=\tau_{d_{2}}$. Let $a\in X$ and $\varepsilon>0$. From (3.14), there is $V$ in $\tau(a)$ such that for every $x,y$ in $V$,
$|d_{1}(x,y)-d_{2}(x,y)|\leq\frac{\min\{\varepsilon,1\}}{2}\,\bigl(d_{1}(x,a)+d_{1}(y,a)\bigr).$ (3.15)
@@
$|d_{2}(x,y)-d_{1}(x,y)|\leq\frac{\varepsilon}{2}\,\bigl(d_{1}(x,a)+d_{1}(y,a)\bigr)\leq\varepsilon\,\bigl(d_{2}(x,a)+d_{2}(y,a)\bigr).$
@@
$\bigl|d_{1}(x,y)-d_{2}(x,y)\bigr|$ $\leq\frac{\varepsilon}{3}\,\bigl(d_{1}(x,a)+d_{1}(y,a)\bigr),$
$\bigl|d_{2}(x,y)-d_{1}(x,y)\bigr|$ $\leq\frac{1}{2}\,\bigl(d_{2}(x,a)+d_{2}(y,a)\bigr),$
@@
$\bigl|d_{2}(x,y)-d_{3}(x,y)\bigr|\leq\frac{\varepsilon}{3}\,\bigl(d_{2}(x,a)+d_{2}(y,a)\bigr).$
@@
$d_{2}(x,a)+d_{2}(y,a)\leq 2\bigl(d_{1}(x,a)+d_{1}(y,a)\bigr).$
Thus,
$\bigl|d_{1}(x,y)-d_{3}(x,y)\bigr|$ $\leq\bigl|d_{1}(x,y)-d_{2}(x,y)\bigr|+\bigl|d_{2}(x,y)-d_{3}(x,y)\bigr|$
$\leq\frac{\varepsilon}{3}\,\bigl(d_{1}(x,a)+d_{1}(y,a)\bigr)+\frac{\varepsilon}{3}\,\bigl(d_{2}(x,a)+d_{2}(y,a)\bigr)$
$\leq\varepsilon\,\bigl(d_{1}(x,a)+d_{1}(y,a)\bigr).\qed$
@@
Let $X$ be a set and $d_{1},d_{2}$ be distances on $X$. If $d_{1}\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}d_{2}$, then $d_{1}\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}d_{2}$.
@@
Suppose that $d_{1}\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}d_{2}$. Let $a\in X$. We use the following upper bound for the expression in (3.13):
@@
As $(x,y)\to(a,a)$ and $x\neq y$, the last expression tends to zero. ∎
Proposition 3.9.
Let $X$ be a set and $d_{1},d_{2}$ be distances on $X$. If $d_{1}\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}d_{2}$, then $d_{1}\cong d_{2}$.
Proof.
Suppose that $d_{1}\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}d_{2}$. Let $a\in X$. We suppose that the limit relation (3.13) holds. Fixing $y=a$ we get the following consequence of (3.13):
$\lim_{x\to a}\frac{|d_{2}(x,a)-d_{1}(x,a)|}{d_{1}(x,a)}=0.$
@@
The next example shows, in particular, that $d_{1}\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}d_{2}$ is not equivalent to $d_{1}\cong d_{2}$.
@@
For two points belonging to the same halfaxis, we have
@@
$\frac{|d_{2}((u,0),(0,u))-d_{1}((u,0),(0,u))|}{d_{1}((u,0),(0,0))+d_{1}((0,u),(0,0))}=\frac{2-\sqrt{2}}{2}.$
@@
Omer Cantor (https://orcid.org/0009-0007-2969-7170) constructed the following example and explained it to us. It shows that $d_{1}\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}d_{2}$ is not equivalent to $d_{1}\stackrel{{\scriptstyle\operatorname{str}}}{{\sim}}d_{2}$.
Example 3.11.
For each parameter $\alpha$ satisfying $0<\alpha\leq 1$, we define $\rho_{\alpha}\colon[-1,1]^{2}\to[0,+\infty)$ by
$\rho_{\alpha}(x,y)\coloneqq\begin{cases}|x-y|,&xy\geq 0;\\ |x+y|+\alpha\bigl(\min\{|x|,|y|\}\bigr)^{2},&xy<0.\end{cases}$
Intuitively, the definition of $\rho_{\alpha}$ means that for each $u$ in $(0,1]$ there is a “teleportation” of cost $\alpha u^{2}$ between $u$ and u-u. Thus, if $0<u<v\leq 1$, then $\rho_{\alpha}(-u,v)$ as obtained as the sum of two terms: the teleportation between u-u and $u$ and the usual walk between $u$ and $v$. It is easy to verify that $\rho_{\alpha}$ is a distance on [1,1][-1,1]. It induces the usual topology on [1,1][-1,1]. Moreover, if $a\neq 0$, then there exists an open neighborhood $V$ of $a$ such that $\rho_{\alpha}(x,y)=|x-y|$ for each $x,y$ in $V$. Consider two different distances from this family, say $\rho_{1/2}$ and $\rho_{1}$. For $a=0$ and $x,y\in[-1,1]$ such that $xy<0$, we have
$\frac{|\rho_{1/2}(x,y)-\rho_{1}(x,y)|}{\rho_{1}(x,0)+\rho_{1}(0,y)}=\frac{\frac{1}{2}\bigl(\min\{|x|,|y|\}\bigr)^{2}}{|x|+|y|}.$
Obviously, this expression tends to 0 as $(x,y)\to(0,0)$. For all other choices of $a,x,y$, the corresponding quotient is equal to 0. Therefore, $\rho_{1/2}\stackrel{{\scriptstyle\operatorname{infin}}}{{\sim}}\rho_{1}$.
On the other hand, for $a=0$, $x_{k}=1/k$, and $y_{k}=-1/k$, we consider the quotient
$\frac{\rho_{1/2}(x_{k},y_{k})}{\rho_{1}(x_{k},y_{k})}=\frac{\frac{1}{2}(1/k)^{2}}{(1/k)^{2}}=\frac{1}{2}.$
As $k\to\infty$, we have $(x_{k},y_{k})\to(0,0)$, but the quotient above does not tend to $1$. Hence, $\rho_{1/2}$ and $\rho_{1}$ are not strongly locally similar.